what is the reciprocal of 9/8?
what is the reciprocal of -2.45?
what is the reciprocal of 3 1/5?
what is the reciprocal of 7/2m?

Answer by checkley71(8403)

RECIPRICAL OF 8/9=9/8
-2.45=1/-2.45
3 1/5=16/5=5/16
7/2M=2M/7